The sum of the ages of Chance and Valerie is 89 years. 12 years ago, Chance's age was 4 times Valerie's age. How old is Chance now?
Answer by JulietG(1812)

C + V = 89 therefore C = 89 - V
C - 12 = 4(V - 12)
C - 12 = 4V - 48
C = 4V - 36
89 - V = 4V - 36
89 = 5V - 36
89 + 36 = 5V
125 = 5V
25 = V
Valerie is 25; therefore Chance is 89-25, or 64.
12 years ago, Valerie would have been 13. Chance would have been 52, which is 4x13